OpenStudy (anonymous):

integral of (sin(3x))^4dx from 0 to pi

OpenStudy (anonymous):

sin^4(3x)=(1/2-1/2cos(6x))^2 1/4-1/2cos(6x)+1/4cos^2(6x) 1/4-1/2cos(6x)+1/4(1/2+1/2cos(12x)) 1/4-1/2cos(6x)+1/8+1/8cos(12x) 3/8-1/2cos(6x)+1/8cos(12x) NOW take the integral of this 3/8x-1/12sin(6x)+1/(96)sin(12x) since sin(pi)=0 because sin(kpi) k being an integer will be 0 3/8pi will be your answer

OpenStudy (anonymous):

NOW check me

OpenStudy (anonymous):

(k) being any whole number

OpenStudy (anonymous):

\[\int\limits_0^{\pi } \left(\text{Sin}[3 x]^4\right) \, dx=\frac{3 \pi }{8} \]